OpenStudy (anonymous):

|2n + 15| = 1. One solution is n = -7.

OpenStudy (anonymous):

Solve the absolute value equation |2n + 15| = 1. One solution is n = -7. What is the other solution? n = 7 none of the above n = -14 n = 8

OpenStudy (anonymous):

It cant be a - number because of | |

OpenStudy (anonymous):

So we have : |2n + 15| = 1 |2n | = 1-15 |2n| = -14 2n=-7but|-7| = +7

OpenStudy (anonymous):

doesn't absolute value mean the answer is positive or negative

OpenStudy (anonymous):

Solve the inequality : -2x - 6 < 4describe the circles and the shading for the graph.

OpenStudy (austinl):

\(|2n + 15| = 1\) \(2n+15=1\) \(-(2n+15)=1\) Solve each of these and they should be your two solutions.

OpenStudy (anonymous):

You should kill the number is before x . ! So divided all on -2 :)

OpenStudy (austinl):

And for the inequality, You would add six to both sides, then divide by negative two and flip the sign.